# Emerging Quadrature Lattices of Kerr Combs

### v1

@@ 4 Methods
4.1 Band model for quadrature lattice induced by a Kerr comb
@@
As shown in Eq. LABEL:eq2 in the main text, 2-FSR Kerr combs generate 1D quadrature lattices on the frequency axis. The lattice geometry allows reducing the complexity of modeling the physics from a general network of quadratures to a 1D lattice geometry. The goal of this section is to explain the physics of the quadrature lattice by deriving an analytical expression for the band model of the quadrature lattice and its Brillouin zone. First, we transform the Hamiltonian in Eq. LABEL:eq2 into the quadrature basis $\vec{R}=(p_{-N},...,p_{N},q_{-N}...,q_{N})^{T}$ where $p_{\mu}=\frac{1}{\sqrt{2}}\left(a_{\mu}^{\dagger}+a_{\mu}\right),q_{\mu}=\frac{i}{\sqrt{2}}\left(a_{\mu}^{\dagger}-a_{\mu}\right)$ and write the quadrature non-Hermitian Hamiltonian $\mathcal{M}$ [46], which satisfies the Heisenberg equation:
@@
We obtain translation symmetry of the lattice by assuming periodic boundary conditions and flat dispersion $\Delta\tilde{\omega}_{\mu}$ in the folded lattice, meaning that $\Delta\tilde{\omega}_{\mu}=\tilde{\delta}_{A}\Theta(\omega-\mu)+\tilde{\delta}_{B}\Theta(\omega+\mu)$ where $\Theta$ is the step function and $\tilde{\delta}_{A(B)}$ are real numbers. We again use the operators $\hat{c}_{\mu},\hat{d}_{\mu}$ and their appropriate quadratures: $\hat{p}_{c,\mu},\hat{q}_{c,\mu},\hat{p}_{d,\mu},\hat{q}_{d,\mu}$ The difference $\tilde{\delta}_{A}-\tilde{\delta}_{B}$ creates an asymmetry in the dispersion $\Delta\tilde{\omega}_{\mu}$ which removes the anti-parity symmetry, thus it functions as a simplified representation of the mismatch $D^{\prime}_{1}$. By using the translation symmetry, we transition to the reciprocal space defined by the good quantum number $k$, obtaining a $4\times 4$ Hamiltonian:
@@
where the variables $g_{A(B),k}=\left(\tilde{\delta}_{A(B)}+2\right)+2\alpha e^{ik}+2\alpha e^{-ik}$ originate from the dispersion, Bragg scattering, and self phase modulation, while $f_{k}=-1-2\alpha e^{ik}-2\alpha e^{-ik}$ are pair generation terms, and $\mathcal{M}=\sum_{k}{\mathcal{M}_{k}+\mathcal{O}(\alpha^{2})}$. The Hamiltonian $\mathcal{M}_{k}$ has pseudo anti-parity symmetry (implying that the eigenvalues of $\mathcal{M}_{k}$ are either purely real or purely imaginary) if $\tilde{\delta}_{A}=\tilde{\delta}_{B}$, which shows why there exist a detuned threshold when $\Delta\tilde{\omega}_{\mu}$ is not symmetric around $\mu=0$.
Figure 6: 2-band model of a quadrature lattice induced by a 2-FSR Kerr comb. a. Illustration of the above threshold comb (red) and below threshold comb (green) in the model, and of the different terms - pair generation and Bragg scattering that couple the different modes. b. The 1D lattice unit cell, in the generalized quadrature basis $\left(\tilde{p}^{\prime}_{\mu},\tilde{q}^{\prime}_{\mu}\right)$. c-e. Different band topologies (different number of exceptional points) of the Brilloin zone of the lattice in b, plot c corresponds to $A_{\pm 2}/A_{0}=0.25,\Delta\tilde{\omega}=2$, plot d corresponds to $A_{2}/A_{0}=0.25,A_{-2}/A_{0}=0.2,\Delta\tilde{\omega}=2$ and plot e to $A_{\pm 2}/A_{0}=0.25,\Delta\tilde{\omega}=0$. f-i. The number of exceptional points for different 2-FSR Kerr combs on different slices of the 3D parameter space. The parameter space is spanned by the two ratios of the sidebands and the overall detuning
@@
Due to the degeneracy, we use only two variables: $\left(p^{\prime}_{k},q^{\prime}_{k}\right)$, allowing us to explore some of the fundamental physics of our lattices. Figure 6 analyzes the 2-band model of Eq. 6. Figure 6(a) shows different non-linear terms and the structure of the comb (red) and sub-comb(green), while Fig. 6(b) presents the folded representation where translation symmetry exists.
@@
Next, we analyze the parameter space of the comb shown in Fig. 6(a) where $\alpha<0.5$. In this parameter sub-space there can exist 0,2 or 4 exceptional points in the Brillouin zone (Fig. 6(c-d)). We map the 3D parameter space for combs (also for combs with $A_{2}\neq A_{-2}$) presenting the different band topologies (Fig. 6(f-i)). We note that having 4 exceptional points is not an inherent maximum; beyond the presented parameter space, more exceptional points can exist. Additionally, the relative phases between the comb tones can influence the number of exceptional points when more than two pumping amplitudes are considered. Therefore, while we provide the basic mechanisms governing such lattices, they can become significantly more complex. The analysis of the 2-band or 4-band models sheds light on the eigen-vector and eigen-value structures. However, the infinite lattice cannot reach threshold by definition and is therefore not suitable for studying multi-mode squeezing phenomena on its own, and must be truncated for that purpose. More details on the truncation and its implication on the results of the 2-band model will be discussed in the next section.
4.2 The finite model and analysis of the robustness of non-detuned quadrature lattice modes
@@
In this section, we continue with the theoretical analysis akin to Methods 4.1  focusing on finite lattices. Similar to structures in physical space, a realistic lattice in the frequency domain is inherently finite. Naturally, dispersion is never completely flat (i.e group velocity dispersion or higher-order terms are never zero), implying that as the lattice extends towards higher and lower frequencies away from the pump, the detuning between the cold cavity dispersion and the quadrature lattice modes increases. This effect reduces the coupling efficiency of Bragg scattering and possibly pair generation, leading to “soft edges” and resulting in a gradual decrease in transport between frequencies. In contrast, edges can also be made abrupt through engineering of mode crossing or other type of frequency dependent defects [20]. These sharp edges can be desirable as they may host confined edge-states in the synthetic frequency dimension [47]. While Kerr resonators can accommodate both options, our focus here is on naturally occurring edges due to non-zero group velocity dispersion.
@@
Furthermore, we will numerically demonstrate that the non-detuned multimode states persist even when dispersion is non-flat and are robust against the breaking of dispersion symmetry around the pump frequency $\mu=0$. Starting from Eq. LABEL:eq2, we assume $\Delta\tilde{\omega}_{\mu}=\omega_{p}+D^{\prime}_{1}\mu+D_{2}\mu^{2}$, where $D^{\prime}_{1}=\Delta\Omega-D_{1}$ and $D_{1}$ represents the FSR. When $D_{2}\neq 0$ and/or $D^{\prime}_{1}\neq 0$ the multimode states localize, allowing us to analyze them using a finite Hamiltonian $\mathcal{M}$. For this analysis, we assume equal quality factors for all resonances $\Gamma=\kappa/2\textbf{I}_{2N\times 2N}$. Since $\Gamma$ is real and proportional to the unit matrix, it only translates the real part of the eigenspectrum of $\mathcal{M}-\Gamma$, and does not influence the imaginary part of the eigenvalues or the peak separation of the squeezed multimode states. Following the conventions in [46] the input-output relation of the system and its Fourier transform are given by:
@@
The linear response matrix $S\left(\omega\right)$ describes in our case how the system amplifies and squeezes the quantum fluctuations in the quadrature lattice and is given by:
@@
The matrix $S\left(\omega\right)$ can be decomposed by the Bloch-Messiah decomposition which leads to:

### v3

@@ 4 Methods
4.1 High-Q 4H-SiC integrated microrings with integrated filters
Our devices are etched from a 550 nm thick 4H-SiC-on-Insulator [56] and are 1930-nm wide, with diameter of 233-um and capped with SiO2 (Fig. 5(a)). Microheaters are patterned via liftoff of 100 nm platinum with a 5 nm titanium adhesion layer. The filters extinction is smaller than 20dB and they are supplemented with off-chip fiber Bragg filters. Coupling to and from the chip is achieved using lensed fibers, with collection efficiency of up to 50%.
Figure 5: Micro-ring resonator: fabrication and dispersion. a. Optical image of the Kerr ring resonator alongside the smaller on-chip filter ring and heaters. b. Normalized transmission over wavelength detuning of a ultra high-Q resonance in our device with intrinsic quality factor of 9 million c. Integrated dispersion of the TE mode of the Kerr ring resonator (blue) and fit to $D_{2}=1.2~{}MHz,D_{1}=152.7$ GHz (red). Dispersion defects are avoided mode crossings with other mode families. d. Cold cavity intrinsic quality factors (Qi - red) and coupling quality factors (Qc - blue) of the resonances in the relevant mode-family.
The ring resonators exhibit intrinsic quality factors of up to 9 million (Fig. 5(b)). The resonators exhibit a $D_{2}/2\pi$ of 1.2MHz, and an FSR of 153 GHz. Additionally, the ring is multimode, causing mode splitting through modal interactions (Fig. 5(c)), which allows a broad range of narrow-bandwidth states to be observed in the devices, such as 2-FSR primary combs, as well as the transition to 1-FSR combs and soliton crystals. The intrinsic quality factors of the mode family used in this work reach 5 million but with average of 1.8 million (Fig. 5(d)). The mode splitting is a source of squeezing reduction due to the mode mismatch at the detector, and the poor output coupling of TM modes. Other factors that reduce the squeezing are the under-coupled to critically coupled resonances and a local oscillator which does not cover 10 cavity modes and mismatch with the oscillating supermode. The heaters modify the mode-splitting properties and the index of refraction, thereby altering the integrated dispersion and the pump detuning at threshold which consequently changes the Kerr comb’s repetition rate $\Delta\Omega$.
4.2 Additional details on the 2-FSR microcombs
In this section we provide additional details and data on the formation process of our 2-FSR microcombs. Figure 6 presents the experimental evolution of the microcomb. The different stages of the comb evolution are tracked by transmission measurements (Fig.6(a)), by an optical spectrum analyzer (OSA) (Fig. 6(b)) and by single photon optical spectrum analyzer (SPOSA) (Fig. 6(c)). In Fig. 6(b) we present the optical spectrum of the comb at different detuning stages. The comb emerges after a first threshold as a 2-FSR primary comb (Fig. 6(b)i-ii) and undergoes a non-oscillating threshold (single peak supermode - as in Fig. 4(d)) to a single 1-FSR phase locked comb (Fig. 6(b)iii). Following the threshold, a second oscillating threshold occurs (double peak - see section 3 in the supplementary material) followed by additional thresholds and RF beat notes in the comb (Fig. 6(b)iv-v). At this point the RF spectrum broadens and becomes chaotic. Finally, the comb abruptly turns to a 2-FSR soliton crystal (Fig. 6(b)vi). The 2-FSR soliton crystal state does not have a long stable 1-FSR soliton state at the end of it. In Fig 6(c) the optical spectrum is measured with an SPOSA. This configuration allows to resolve the above and below threshold light in the same measurement sequence due to the high dynamical range. Figure 6(c) shows how some of the below threshold light is generated far away from the pump (around 1520 and 1570nm), but also in the center modes (between 1540nm-1550nm), which is where we primarily measure.
Figure 6: The experimental optical spectrum of the Kerr microcomb. a. Transmission as a function of pump detuning. The different colored lowercase Roman numerals correspond to the different stages of the microcomb. b. The spectrum of the different combs specified by lowercase Roman numerals in (a). i and ii correspond to the primary 2-FSR comb which is the main focus of this work, while iii is a 1-FSR comb, and iv-v are chaotic states. Finally vi is a 2-FSR stable soliton crystal state. c Single-photon optical spectrum analyzer (SPOSA) measurement of photon populations in above threshold (red) and below threshold (green) modes of the 2-FSR state.
4.3 Additional data on LLE simulation
In order to support the theoretical conclusions and experimental observations, we preformed LLE simulations of a 2-FSR Kerr comb. The LLE simulates the classical comb dynamics [40], and allows us to better understand the relation of oscillating and non-oscillating squeezed structures, on the formation of secondary combs.
We start by simulating a dispersion with a mode crossing which means that we locally detune in frequency resonance near the pump. The detuning yields a comb similar to that of our experiment. The resonance frequency of the mode crossing changes in response to heating. For the same reason, the detuning of the resonances close to the mode crossing also changes as a function of power in the cavity (i.e it is changing during the evolution of the comb). However, since the effects we describe do not require this behavior, we can observe the main features under constant dispersion. We use experimentally extracted parameters of $D_{2}/2\pi=1.2MHz$ and extract the dispersion perturbation (from the mode crossing) parameters from the measured dispersion (Methods 4.1). For simplicity, we only perturbe the modes near the pump modes $-2,-1,0,1,2$. For quality factors of the rest of the resonances, we use a constant value of 1.83 million which is the measured average value, and we reduce the intrinsic quality factors at modes distant 17 FSR away from the pump to avoid an opo threshold there as in the measured device. The normalized pump power is $f=5.93$ where $f^{2}=8g_{0}\kappa_{c}P_{in}/\kappa^{3}\hbar\omega_{0}$ and $P_{in}$ is the input power, $\kappa_{c}$ is the out coupling power loss, $\omega_{0}$ is the pump angular frequency, and $\hbar$ is the reduced Planck’s constant.
Importantly, the formation of our 2-FSR comb is not a direct result of the anomalous dispersion, but it is formed locally due to the mode crossing in the dispersion (see Methods  4.1). This means that the initial comb formation can be modified by adjusting the frequency of cavity resonances $-2,0,2$. The pump detuning value of the threshold to a 1-FSR comb depends strongly on the dispersion of the odd-numbered modes and the perturbation of modes $-1,1$. By adjusting these parameters, we are able to simulate a comb evolution that is qualitatively similar to what we observe in the experiment. Hence tuning them reduces the oscillations of the supermode until it reaches a region with non-oscillating supermode, similarly to the effect of changing the voltage on the microheaters in the experiment.
Figure 7(a) shows the comb evolution in the simulation, which is similar to our experiment (Fig. 6). The process begins with a 2-FSR primary comb which transitions to a 1-FSR secondary comb. Figure 7(b)i demonstrates the transitions from one rolling Turing pattern (2-FSR) to another (1-FSR). Since this transition is non-oscillating, the 1-FSR comb is a single phase locked comb and does not have sub combs (shown experimentally in section 3 in the Supplementary Material).
Figure 7: Intracavity field in real space for different regimes in the comb formation - LLE with perturbed dispersion - non oscillating threshold. a. Transmission as a function of continuous pump detuning (toward longer wavelengths). This plot is the same as in Fig. 1(b) but the shaded areas mark different regimes plotted in (b). b. Intracavity field in real space for the transition between the (i): primary comb (2-FSR) and the secondary comb (1-FSR) corresponding to the shaded red region in (a). (ii): the transition from the secondary comb (1-FSR) to a second oscillating 1-FSR comb which oscillates (shaded green region). Experimental evidance of this transition can be found in section 3 of Supplementary Material. (iii): The region of a 2-FSR soliton crystal.
The next (third) transition involves the threshold crossing of an oscillating 1-FSR supermode, which produces an RF beat note with the existing 1-FSR comb. The oscillations in this state are evident in both the transmission and the intra-cavity field (Fig. 7(b)ii). Finally the system transitions to a 2-FSR soliton crystal (Fig. 7(b)iii). To better understand the unique non-oscillating transition from a 2-FSR comb to a single phase locked 1-FSR comb, we explore the comb state at a pump detuning close to this threshold at 0.115 GHz.
Figure 8: Examination of the simulated threshold between a 2-FSR comb to a 1-FSR comb at 0.115 GHz detuning. a. Spectrum of the 2-FSR primary Kerr comb (red) and light in odd-numbered modes (green). b. Intracavity field in real space (rolling Turing) c. $g^{(2)}$ correlations of the below threshold light. d. Eigen-spectrum of the comb showing that a non-oscillating supermode crosses threshold, and that the spectrum also has oscillating supermodes with less parametric gain. e. The squeezing RF spectrum, showing the single peak which correspond with the non-oscillating supermode that crosses threshold in d. f. The spectrum of the most squeezed state in the basis of cavity modes.
Figure  8(a-b) shows the comb from Fig. 7 for the detuning of 0.115 GHz in both the frequency (mode) space and real physical space respectively. Using the obtained comb amplitudes and the hot-cavity dispersion, we calculate the $g^{(2)}$ correlations of each cavity-mode $\mu$ with each cavity-mode $\nu$, which resembles the experimental $g^{(2)}$ in Fig. 2(c). The oscillating or non-oscillating nature of the supermode is not visible from Fig. 2(c) and requires the temporal lineshapes. Next, we calculate the quadrature Hamiltonian $\mathcal{M}$ associated with the 2-FSR comb at 0.115GHz. The eigenspectrum is plotted in Fig. 8(d), showing that the non-oscillating state (purely real eigenvalue) is approaching threshold. Explicitly, this is evidant from the fact that the supermode with the highest gain (blue dot closest to zero) has zero oscillations (red dot at zero). Correspondingly, the squeezing structure is a single peak (Fig. 8(e)), centered around the pump (Fig. 8(f)). When this supermode crosses threshold it does generate RF oscillations.
4.4 Additional details on the experimental setup
Here, we provide a more detailed description of the experimental setup. Figure  9(a) illustrates the fiber based experimental setup around the microchip. We utilize two tunable CW telecom lasers named CW laser 1 (Velocity, 1520-1570 nm, TLB-6728) and CW laser 2 (Toptica CTL1500: 1460-1570nm). The lasers are split into two channels. The first channel (upper channel in Fig. 9(b)) receives light from laser 1 and after amplification through an EDFA excites the Kerr resonator. The Kerr resonator outputs two channels: the first carries the quantum signal with the filtered Kerr comb directly to the homodyne unit, while the second channel contains the Kerr comb passing through the filter ring to a photodiode.
The second channel (bottom in Fig. 9(b)), receives light either from CW laser 1 or CW laser 2 or both. CW laser 1 is used for measuring modes -1 and 1, while CW laser 2 allows measuring other modes and is used for noise spectroscopy in Fig.3(c).
4.5 Band model for the lattice dynamics induced by a Kerr comb
@@
As shown in Eq. LABEL:eq2 in the main text, 2-FSR Kerr combs generate 1D lattice dynamics on the frequency axis. The lattice geometry allows reducing the complexity of modeling the physics from a general network of quadratures to a 1D lattice geometry. The goal of this section is to analyze the lattice dynamics by deriving an analytical expression for the band model of the lattice and its Brillouin zone. First, we transform the Hamiltonian in Eq. LABEL:eq2 into the quadrature basis $\vec{R}=(p_{-N},...,p_{N},q_{-N}...,q_{N})^{T}$ where $p_{\mu}=\frac{1}{\sqrt{2}}\left(a_{\mu}^{\dagger}+a_{\mu}\right),q_{\mu}=\frac{i}{\sqrt{2}}\left(a_{\mu}^{\dagger}-a_{\mu}\right)$ and write the quadrature non-Hermitian Hamiltonian $\mathcal{M}$ [57], which satisfies the Heisenberg equation:
@@
We obtain translation symmetry of the lattice by assuming periodic boundary conditions and flat dispersion $\Delta\tilde{\omega}_{\mu}$ in the folded lattice, meaning that $\Delta\tilde{\omega}_{\mu}=\tilde{\delta}_{A}\Theta(\omega-\mu)+\tilde{\delta}_{B}\Theta(\omega+\mu)$ where $\Theta$ is the step function and $\tilde{\delta}_{A(B)}$ are real numbers. We again use the operators $\hat{c}_{\mu},\hat{d}_{\mu}$ and their appropriate quadratures: $\hat{p}_{c,\mu},\hat{q}_{c,\mu},\hat{p}_{d,\mu},\hat{q}_{d,\mu}$. The difference $\tilde{\delta}_{A}-\tilde{\delta}_{B}$ creates an asymmetry in the dispersion $\Delta\tilde{\omega}_{\mu}$ which removes the anti-parity symmetry, thus it functions as a simplified representation of the mismatch between the comb repetition rate and dispersion $D^{\prime}_{1}$. By using the translation symmetry, we transition to the reciprocal space defined by the good quantum number $k$, obtaining a $4\times 4$ Hamiltonian:
@@
where the variables $g_{A(B),k}=\left(\tilde{\delta}_{A(B)}+2\right)+2\alpha e^{ik}+2\alpha e^{-ik}$ originate from the dispersion, Bragg scattering, and self phase modulation, while $f_{k}=-1-2\alpha e^{ik}-2\alpha e^{-ik}$ are pair generation terms, and $\mathcal{M}=\sum_{k}{\mathcal{M}_{k}+\mathcal{O}(\alpha^{2})}$. The Hamiltonian $\mathcal{M}_{k}$ has pseudo anti-parity symmetry (implying that the eigenvalues of $\mathcal{M}_{k}$ are either purely real or purely imaginary) if $\tilde{\delta}_{A}=\tilde{\delta}_{B}$, which shows why there exist an oscillating threshold when $\Delta\tilde{\omega}_{\mu}$ is not symmetric around $\mu=0$.
Figure 10: 2-band model of a quadrature lattice induced by a 2-FSR Kerr comb. a. Illustration of the above threshold comb (red) and below threshold comb (green) in the model, and of the different terms - pair generation and Bragg scattering that couple the different modes. b. The 1D lattice unit cell, in the generalized quadrature basis $\left(\tilde{p}^{\prime}_{\mu},\tilde{q}^{\prime}_{\mu}\right)$. c. Different band topologies (different number of exceptional points) of the Brilloin zone of the lattice in b, top plot corresponds to $A_{\pm 2}/A_{0}=0.25,\Delta\tilde{\omega}=2$, center plot corresponds to $A_{2}/A_{0}=0.25,A_{-2}/A_{0}=0.2,\Delta\tilde{\omega}=2$ and bottom plot to $A_{\pm 2}/A_{0}=0.25,\Delta\tilde{\omega}=0$. d. The number of exceptional points for different 2-FSR Kerr combs on different slices of the 3D parameter space (i-iii). The parameter space is spanned by the two ratios of the sidebands and the overall detuning (bottom right).
@@
Due to the degeneracy, we use only two variables: $\left(p^{\prime}_{k},q^{\prime}_{k}\right)$, allowing us to explore some of the fundamental physics of our lattices. Figure 10 analyzes the 2-band model of Eq. 6. Figure 10(a) shows different non-linear terms and the structure of the comb (red) and sub-comb(green), while Fig. 10(b) presents the folded representation where translation symmetry exists. We stress that in the analysis shown in Fig. 10 and subsequent simulations we take into acount also next-nearest neighbor terms. These terms preserve the anti-parity time symmetry in the model. Additionally, in the regime of $\alpha<0.5$ the effect of the next nearest neighbor terms on the phase maps were found to be marginal as expected.
Next, we analyze the parameter space of the comb shown in Fig. 10(a) where $\alpha<0.5$. In our experiment $\alpha$ reaches 0.355, which is well within this region. We do not address the region of $\alpha>0.5$ or very broad combs in this work, as they require separate study and are beyond the scope of this paper. Regarding the detuning parameter $\tilde{\Delta\omega}$, this parameter varies in the experiment as a function of the pump detuning and the voltage in the microheater. Therefore, we keep it as a free parameter and sweep over it in the widest relevant range. In this parameter sub-space there can exist 0,2 or 4 exceptional points in the Brillouin zone (Fig. 10(c)). We map the 3D parameter space for combs (also for combs with $A_{2}\neq A_{-2}$) presenting the different band topologies (Fig. 10(d)).
@@
From the phase maps in Fig. 10(d), our experimental parameter $\alpha$ (which is also close to the LLE simulation) is expected to have a broad detuning ($\Delta\tilde{\omega}$) region where non-oscillatory states should exist. Additionally, non-oscillatory supermodes are more likely to appear with higher $\alpha$ values. We note that having 4 exceptional points is not an inherent maximum; beyond the presented parameter space, more exceptional points can exist. Additionally, the relative phases between the comb tones can influence the number of exceptional points when more than two pumping amplitudes are considered. Therefore, while we provide the basic mechanisms governing such lattices, they can become significantly more complex. The analysis of the 2-band or 4-band models sheds light on the eigenvector and eigenvalue structures. However, the infinite lattice cannot reach threshold by definition and is therefore not suitable for studying multi-mode squeezing phenomena on its own, and must be truncated for that purpose. More details on the truncation and its implication on the results of the 2-band model will be discussed in the next section.
4.6 The finite model analysis
@@
In this section, we continue with the theoretical analysis akin to Methods 4.5, focusing on finite lattice dynamics. Similar to structures in physical space, a realistic lattice in the frequency domain is inherently finite. Naturally, dispersion is never completely flat (i.e group velocity dispersion or higher-order terms are never zero), implying that as the lattice extends towards higher and lower frequencies away from the pump, the detuning between the cavity dispersion and the repetition rate increases. This effect reduces the coupling efficiency of Bragg scattering and possibly pair generation, leading to “soft edges” and resulting in a gradual decrease in transport between frequencies. In contrast, edges can also be made abrupt through engineering of mode crossing or other type of frequency dependent defects [47]. These sharp edges can be desirable as they may host confined edge-states in the synthetic frequency dimension [44]. While Kerr resonators can accommodate both options, our focus here is on naturally occurring edges due to non-zero group velocity dispersion. Furthermore, we will numerically demonstrate that the non-oscillating multimode states persist even when dispersion is not completely symmetric and in the finite regime
@@
We start from Eq. LABEL:eq2, and assume $\Delta\tilde{\omega}_{\mu}=\omega_{p}+D^{\prime}_{1}\mu+D_{2}\mu^{2}$, where $D^{\prime}_{1}=\Delta\Omega-D_{1}$ and $D_{1}$ represents the FSR. When either $D_{2}\neq 0$ or $D^{\prime}_{1}\neq 0$ the multimode states become finite in their spectrum, allowing us to analyze them using a finite Hamiltonian $\mathcal{M}$. For this analysis, we assume equal quality factors for all resonances $\Gamma=(\kappa/2)\textbf{I}_{2N\times 2N}$, where, for simplicity, $\kappa$ is the uniform intensity loss rate. Since $\Gamma$ is real and proportional to the unit matrix, it only translates the real part of the eigenspectrum of $\mathcal{M}-\Gamma$, and does not influence the imaginary part of the eigenvalues or the peak separation of the squeezed multimode states. Following the conventions in [57], the input-output relation of the system and its Fourier transform are given by:
@@
The linear response matrix $S\left(\omega\right)$ describes in our case how the system amplifies and squeezes the quantum fluctuations and is given by:
@@
To describe the quantum vacuum we solve the frequency dependent linear response. [57]. We obtain the Morphing supermodes which describe the dynamical nature of the light. The symplectic form of $\mathcal{M}$ results in a decomposition that is a smooth function of the real parameter $\omega$. Thus, when we apply the Bloch-Massiach decomposition to $\mathcal{M}$ we obtain: